18.034 PROBLEM SET 10

Due May 5 in class. No lates will be accepted. Discussion is encouraged, with two caveats: (a) write up your solutions by yourself, and (b) give credit when others came up with ideas (you won’t be penalized for this). Give explanations, not just answers. References are to Boyce and DiPrima; the answers to problems from the book are in the back of the book.

1. Problem 5.4.2, p. 253.

2. Problem 5.4.21, p. 254.

3. Problem 5.4.24, p. 254.

4. Problem 5.5.17, p. 260.

5. Problem 5.5.23, p. 261.

6. Problem 5.5.25, p. 261.

7. Problem 5.6.4, p. 266.

8. Problem 5.6.11, p. 266.

Date: April 28, 2000.